in week 1, 80 people attended a basketball game.
in week 2, 120 people attended a basketball game.
by what percent did the attendance increase from week 1 to week 2?
120-80=40 - correct.
120.00-.33=119.67 - not exactly sure what you were doing here.

Answer by Theo(13342)

week 1 = 80
week 2 = 120
120 - 80 = 40
40 / 80 = 1/2 = .5 = 50%

the increase was 40.
the original amount was 80.
the percent increase was 40/80 = 4/8 = 1/2 = .5 * 100 = 50%.

the increase from the original amount has to be divided by the original amount to get the percent increase from the original amount.